Back to the NIKA2 main page

Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote. First Pool week: Nicolas Ponthieu, Alexandre Beelen, Juan Macias, Laurence Perotto, Florian Kéruzoré, Stavroula Katsioli, Hélène Roussel, Florian Ruppin. A new sweep was performed.

Oct 20th, Tuesday

High winds with gusts up to 40m/s

Oct 21th, Wednesday

With hail!

No observation could be performed so far.

Oct 22th, Thursday

It's even raining in Granada. No observations.

Oct 23rd, Friday

Morning shift:

We can start the observations, the opacity is excellent and the weather is much better, but we can't start science because we lack a beammap to reduce data. NP, AB and BL perform observation of Venus with NP reducing the beammaps successfully. Grenoble is warned about the presence of the beammaps, but the quicklook with PIIC won't be ready before Monday. The IDL pipeline is used instead!

We let the weather slowly stabilize until we can finally start observations after the sunset. Opacity throughout the evening: ~0.2

LPSZ starts science for this pool (ACT-CL-J0223)! ...

Oct 23th, Friday

03h00 - 09h00; Stavroula and Hélène: After the end of the project 199-16, we observed IMEGIN targets NGC628, NGC925 and NGC2841. Opacity had values under 0.25 and the conditions were stable. Box S in array 3 has been missing for all observations so far (not clear whether it's been lost in the new sweep, or if it can be recovered with a new beammap).

During the morning shift it was observed the project 074-20 and 077-20.

Source of 077-20 is 5 degrees from the sun and while it did not hurt the science observations for a while, it ended up dramatically changing the beam parameters for the worse. So we let the antenna relax in front of M16 (029-20)

Oct 24th, Saturday

03h00 - 09h00; Stavroula and Hélène: After the end of the project 199-16, we observed IMEGIN targets NGC628, NGC925, NGC2841, NGC2976, and 037-20 after that. Opacity had values below 0.20.

09h00 - 15h00; Florian Ruppin: Started the shift with a skydip. After a succession of pointing/focus, started to observe IDCSJ1426 (project 077-20) for 2 hours. Switched to project 074-20 and observed RXJ1347 for 1 hour.

21h00 - 03h00; Laurence: Opacity starts around 0.3 and continuously drops during the shift. We started with observations of a source at high elevation for 039-20 at tau~0.25 and continued observing until the source was too high. After a focus-pointing sequence on Uranus, we did a calib_1scan, which will be useful for the absolute calibration. Then, around 00:00 UT, we moved on to LP-SZ. We observed PSZ2G155 until it was at an elevation as low as 40 degrees and then, we picked ACT-CL-J0240 just after the transit at el > 50deg

Oct 25th, Sunday

03h00 - 09h00; Stavroula and Hélène: Good weather, opacities around 0.2. Observed N628 (briefly), N925, N2841 and N2976 for project 160-16. Should keep integrating on these targets until desired sensitivity achieved. A calib_1scan on Uranus and skydip

09h00 - 15h00; Florian Ruppin: Did exactly the same program as yesterday with more efficiency (less pointing/focus at the beginning of the shift) as the weather conditions were very stable.

21h00 - 24h00; Florian K.: Started the run with observations of secondary calibrators (MWC349, NGC7027, CRL2688). Continued with observations of 039-20 (HD15745).

Oct 26th, Monday

03h00 - 09h00; Stavroula and Hélène: Bad weather, no observations.

09h00 - 15h00; Florian Ruppin: Good weather is back. Started the observations with a skydip and a lot of pointing/focus as the beam was very unstable due to de-icing. Started to observe RXJ1347 (project 074-20) for 3 hours as soon as the beam properties came back to average values.

00h00 - 03h00; Florian K.: Had to stop because of the strong opacity jump.

15h00 - 21h00; Ángel: Opacity about 0.15-0.20. Observed 6 reps. of G82 (007-20) and 1 rep. of HD143894 (039-20).

21h00 - 24h00; Florian K.: Started the run with 007-20 (one execution of observe_G82 to complete the target). Ran the secondary calibrators (MWC349, NGC7027, CRL2688). Observed 199-16 (two repetitions of observe_PSZ2G155 to complete the target). Made a beammap on Uranus at tau_225=0.18, el=60deg (20201026s217).

Oct 27th, Tuesday

00h00 - 03h00; Florian K.: Together Ángel performed a pointing session from 23:15 to 01:15 (UT) - 61 raw sources. Went back to 199-16 for a repetition of observe_ACTJ0240.

03h00 - 09h00; Stavroula and Hélène: Opacities between ~ 0.3 and 0.5. Observed IC443G for project 037-20. Toward the end of the shift, opacities are quickly rising and conditions degrade. Note in TAPAS for 037-20 that the latest scans may be unusable and should be checked. Desired total integration time not yet achieved. A calib_1scan on Uranus and skydip were carried out.

08h00 - 09h30; Matthieu: Stopped at 09h00 because of high opacity. Maintenance works since then.

16h00 - 21h00; Ángel: High opacity (0.7<tau<1, strong instability). Switching to EMIR at ~18h for stability monitoring. Backup observations with EMIR (project 073-20) with slow opacity decreasing. Resuming with NIKA2 at ~21h00.

21h00 - 24h00; Miren, Jean-François & Juan: Pretty bad weather. Opacity at about 0.6. We observed 029-20 following Bilal instructions.

DailyReportsNika2Pool201020 (last edited 2020-11-05 10:32:23 by NikaBolometer)